OpenStudy (anonymous):

Please help factor the trinomial v^2-11v+28

OpenStudy (anonymous):

(v-7)(v-4) -7-4 = -11 -7*-4 = 28

OpenStudy (anonymous):

thank you I am struggling through this :)

OpenStudy (anonymous):

yw